OpenStudy (anonymous):

a certain number decreased by 7 multiplied by the sum of the number and 7. the product is 51. find the number.

jhonyy9 (jhonyy9):

51=17x3

OpenStudy (anonymous):

(x-7)(x+7)= 51 x^2 - 49 = 51 x^2 = 100 x = 10, -10